Solve

xy – 5y + 4x1/2y = 0.
Answer #1

U sod:- xy_5y+482y =o Divide with & on both sides y. 5 y y =0 + 3 (22)% ED compare equo with y+p(a)y+Q(2)y=1 Here, plK+2 -1 E an (n) =D 00 KADA \+2 an(n) an (a) + 4 =0 nco no S] [CMQ) ( ^_2-1) an ouer as 3. [(-2)(*19-6))on ca . [(me)(m2-1) J7 a,= D |«+ 9-) Camare Coefficent of (a) both sides ha 3, 4, 5, - 2) (+ 2) (+ 2 - ) n + 4-а, -2 = 0; 400-3 а ) an = ( 40 ) (Take kao 3 а) ус - Г. а. — 4.aь 3/ - Lx)4 (-21/4 ) 120, (-я) (36)-А) t To4 - 6 (-4op) »» У (x) -t - (1 (3) 1 .... [1-3 см, ..

answered by: ANURANJAN SARSAM